Grade 11Algebra

find the values of b for which the equation 2log1/25(bx+28) = -log5(12-4x-x2) has only one solution.

ankit  agrawal
14 Years agoGrade 11

3 Answers

Chetan Mandayam Nayakar
14 Years ago

2log1/25(bx+28) = -log5(12-4x-x2), log(1/25)(bx+28)2=log5(12-4x-x2)-1

(12-4x-x2)-1=bx+28, (bx+28)(12-4x-x2)=1

 

Swapnil Saxena
14 Years ago

= 2 log1/25 (bx+28) = -log5(12-4x-x2)

= -2 log25 (bx+28) = -log5(12-4x-x2)

= - log5(bx+28) = -log5(12-4x-x2)

= log5(bx+28) = log5(12-4x-x2)

= bx+28 = 12-4x-x2

= x2+ (4+b)x +16

If the quad has only one sol. then D=0

(4+b)2-4(16) =  b2 + 8b - 48=0

(b+12)(b-4)=0

b=12,4


saket shrivastava
14 Years ago

method as above ans is -12, 4 for x belongs (-6,2)